Employees of a discount appliance store receive an additional 20% off of the lowest price on an item. If an employee purchases a dishwasher during a 15% off sale, how much will he pay if the dishwasher originally cost $450?
Answer by KnightOwlTutor(293)

convert % into decimal
20%=20/100=.2
15%=15/100=.15
Sale price of the dishwasher =$450-$450(.15)=$450-$67.5=$382.50
additional discount of .2(sale price of item)=$ off
Cost of diswasher to the employee = $382.5 -$382.50(.2)=$306